Assuning that the inflation rate frum 2015 through 2017 is 6% per year, what would be the real value or a $50.000 income in each of the three years? (Show all work) What does the real value for 2016 mcan? l.
Answer #1

Q. Assuming inflation rate is 6% per year, real values will be:

If one have $50,000 in the year 2014

Real value of $50,000 in 2015 = 50000*(1-0.06) = 50000*0.94 = $47,000

Real value of $50,000 in 2016 = 50000*(1-0.06)^2 = 50000*0.94*0.94 = $44,180

Real value of $50,000 in 2017 = 50000*(1-0.06)^3 = 50000*0.94 *0.94*0.94 = $41529

Real value for the year 2016 which is equal to $44,180 means that the value of money has degraded fron $50,000 in year 2014. Due to inflation value of money as well as its purchasing power decreases. Thus real value of money indicate the purchasing power of money considering that inflation has increased the prices of the products.